# DynoSAM: Open-Source Smoothing and Mapping Framework for Dynamic SLAM

### v2

@@ VI-B1 Feature Detection & Tracking
Static Feature Tracking. At each time-step, a sparse set of detected static keypoints $\mathcal{S}_{\text{2D},k}$ are tracked across consecutive frames using optical flow. Our framework offers options to select different feature extraction algorithms depending on the use case; for our experiments we used Shi-Tomasi corners [57] for indoor environments and ORB [58] features for outdoor environments. By default, our front-end uses the Lucas-Kanade tracker [59] to generate feature correspondences. The input optical flow $I^{F}_{k}$can also be used when it is reliable. Once tracked, our front-end performs geometric verification to discard poor correspondences.
@@
To improve robust tracking, we retain only relevant features by applying Adaptive Non-Maximal Suppression (ANMS) [60] to the detected keypoints. This procedure retains only the most informative and reliable features, while simultaneously promoting a spatially uniform distribution of features across the image. ANMS achieves this by culling uninformative features whilst ensuring that a minimum number of features remain ($800$ in our experiments). Our front-end will detect new keypoints on the static structure if the number of inlier tracks falls below this minimum threshold. This adaptive mechanism ensures that a consistently high number of well-distributed features are tracked across successive frames, thereby enhancing tracking robustness and accuracy.
@@
Dynamic Feature Tracking. For dynamic objects, we seek to track a dense set of features for each object, $\mathcal{D}_{\text{2D},k}$. Dense tracking is important to achieve good coverage over the entire observable object for robust motion estimation [51]. For each dynamic measurement we maintain an associated object-level label $j$, such that features can be associated to the same object between frames.
@@
The instance segmentation mask $I^{S}_{k}$is used to retrieve the object label $j$ per pixel. While this mask can be generated from any standard instance segmentation network, many of these networks do not guarantee that the instance labels will be temporally consistent. Before tracking dynamic keypoints, object masks are tracked using [61]. We then remap the original instance labels to a temporally consistent label $j$ per object. If the user has their own instance segmentation and tracking method, they can choose to bypass this algorithm. In that case, the user only needs to ensure that each pixel represents the unique tracking label $j$.
@@
Although the instance segmentation mask $I^{S}_{k}$is able to separate objects from the background, it cannot distinguish between static and moving objects. To address this, we use the method from [51] to identify moving objects based on scene flow, allowing us to focus on tracking features only on the dynamic objects. This module can be easily modified to integrate learning-based approaches to identify moving objects in the scene.
@@
Once the moving objects in the image are identified and tracked, dense features are extracted by sampling keypoints uniformly within each tracked object mask. The optical flow image $I^{F}_{k}$is then used directly to find the correspondences between frames. Our tracking algorithm algorithm ensures that, where possible, a consistent and significant number of features ($800$ in our experiments) on each object are maintained and tracked. Similarly to static points, only inliers from the previous frame ae used.
VI-B2 Initial Camera Pose Estimation
@@
At each time-step, static measurements $\mathcal{S}_{k}$ are used to estimate the initial camera pose. As shown in the motion estimation module in Fig. 5, the first step is the PnP algorithm [62] which estimates $\prescript{W}{}{\mathbf{X}}_{k}$ by minimizing the re-projection error:
@@
between the tracked static keypoints in the current frame at time-step $k$ and the local map constructed from the previous frame. RANSAC verification and outlier rejection is performed using the implementation as provided by OpenGV [63] to obtain robust initial estimation of the camera pose. This initial camera pose is further refined through joint optical flow and pose optimization method detailed in Sec. VI-B4.
VI-B3 Initial Object Motion Estimation
The initial object motion $\prescript{W}{k-1}{\mathbf{H}}_{k}\in\mathcal{H}_{k}$ is estimated following the method of Zhang et al[15]:
@@ VI-B3 Initial Object Motion Estimation
where $\prescript{W}{k-1}{\mathbf{G}}_{k}=\prescript{W}{}{\mathbf{X}}_{k}^{-1}\>\prescript{W}{k-1}{\mathbf{H}}_{k}$. Since Eq. (26) is in the same form as Eq. (24), we can solve for $\prescript{W}{k-1}{\mathbf{G}}_{k}$ in the same fashion, and the object motion:
@@
can be directly recovered. As before, RANSAC is used to detect outliers and update the inlier tracks in the current frame. Building on [15], we enhance the initial object motion estimation and refine the inlier tracks through two additional steps, detailed in Sec. VI-B4 and Sec. VI-B5, respectively. The method proposed in  Sec. VI-B4 was used in [15] to improve the estimation and tracking. With the addition of  Sec. VI-B5 our method augments this process to further improve the estimation and identify outliers. At each stage of the refinement process, the set of inlier tracks is updated, significantly improving the robustness of the estimation. The three-steps outlined assumes that the motion of each object is independent, allowing DynoSAM to parallelize the object motion estimation process by handling each object instance in a separate thread.
VI-B4 Joint Optical-Flow Refinement
Following the work of [15, 51], the optical flow is jointly refined with the camera pose and object motions. This step is important to ensure robust and accurate feature tracking on both static and dynamic features. The measured flow and the associated $\mathrm{SE}(3)$ transformation are jointly refined by reformulating the re-projection errors from Eq. (26) and Eq. (24) in terms of the measured flow:
$\mathbf{r}_{f,H}$ $=\mathbf{z}_{\text{2D},k-1}+\mathbf{f}_{k-1,k}-\pi(\prescript{W}{k-1}{\mathbf{G}}_{k}\>\prescript{W}{}{\mathbf{m}}_{k-1}^{i})$ (27)
$\mathbf{r}_{f,X}$ $=\mathbf{z}_{\text{2D},k-1}+\mathbf{f}_{k-1,k}-\pi(\prescript{W}{}{\mathbf{X}}_{k}^{-1}\>\prescript{W}{}{\mathbf{m}}^{i})$
@@ VI-B4 Joint Optical-Flow Refinement
where $\theta$ represents either $\prescript{W}{}{\mathbf{X}}_{k}$ camera pose or $\prescript{W}{k-1}{\mathbf{G}}_{k}$ in the case of object motion estimation. Depending on $\theta$, $\mathbf{r}_{f,\theta}$ is either $\mathbf{r}_{f,X}$ or $\mathbf{r}_{f,H}$, as in Eq. (27). The covariance matrix $\Sigma_{0}\in\mathbb{R}^{{2\times 2}}$ associated with the flow prior $\mathbf{r}_{0}(f)$ is diagonal and associated with the measured optical-flow. As each object moves independently, we construct and solve this optimization problem in parallel per object. Through this joint optimization, we refine temporal point associations improving feature tracking and pose estimation. During optimization, we enhance robustness by eliminating additional outliers, specifically those exhibiting large re-projection errors.
VI-B5 Object Motion Refinement
The object motion estimates can be further refined by employing the rigid-body motion model defined in Eq. (17) to formulate an additional nonlinear least-square optimization problem that directly estimates the object motion $\prescript{W}{k-1}{\mathbf{H}}_{k}$. In contrast to the preceding step, which uses a 2D pixel error, this residual function is based on 3D point errors. Empirically, we have observed that this 3D-based approach significantly enhances estimation accuracy and removes additional outliers not previously identified.
Using the motion estimate from the previous steps as an initial estimate, the motion only refinement graph shown in Fig. 6 is build per object and defines the nonlinear least-squares problem:
@@ VI-B5 Object Motion Refinement
The Levenberg–Marquardt solver is used to obtain the final object motion estimate $\prescript{W}{k-1}{\mathbf{H}}_{k}$. Based on Eq. (24), Eq. (29b) forms projection factors with an associated covariance matrix $\Sigma_{2D,R}\in\mathbb{R}^{2\times 2}$, connecting dynamic points and the observing camera poses at time-step $k-1$ and $k$. Eq. (29c) describes the 3D motion residual of Eq. (17) and connects the common motion with each tracked landmark. $\Sigma_{3D,H}\in\mathbb{R}^{3\times 3}$ is the associated covariance matrix. Eq. (29a) represents prior factors on the camera poses. Since this refinement step focuses exclusively on improving object motion estimation, and not camera pose, a strong prior on the camera poses are used during optimization. This is achieved by assigning the covariance matrix $\Sigma_{0}\in\mathbb{R}^{6\times 6}$ with very small values ($\sigma=0.0001$ in our experiments).
@@
Figure 6: Factor graph representing the motion only refinement graph. For each object $j\in\mathcal{J}_{k}$, the motion is refined by minimizing the 3D motion residual using measurements of points tracked between $k-1$ and $k$. In our example we show the dynamic points in orange, the motion in green, the 3D motion residual in purple and the re-projection factors in white. We only show two sets of points $i$ and $i+n$, while in reality there would be $n$ motion factors.
VI-C Back-end
@@
The back-end fuses static and dynamic measurements to jointly estimate the trajectory of the camera and each object as well as a global map of static and dynamic points. Our back-end allows different Dynamic SLAM formulations to be selected and used for estimation. Full-batch and sliding window optimization methods are provided.
As discussed in [18], the optimization problem for Dynamic SLAM can be formulated in various ways, employing different sets of variables and objective functions. The choice of formulation plays a crucial role in determining the system’s accuracy. In this work, we present and evaluate two world-centric estimation approaches, described in Sec. V, as integral components of DynoSAM.
@@
To enable the seamless use of the DynoSAM pipeline for downstream tasks, the system’s output must be clearly defined and consistently maintained, regardless of the specific formulation used to solve the estimation problem. As previously discussed, formulations can represent objects using different state variables (poses or motions as shown in this work) and in different frames of reference (locally as in [21] or in the camera frame [17]). Therefore, a post-estimation processing step is required to convert all estimated states into a common representation to ensure that results can be directly compared, evaluated using consistent metrics, and readily used by downstream applications. Our framework facilitates this by clearly defining a set of outputs:
$\mathcal{O}_{k}=[^{w}X_{k},\prescript{W}{k-1}{\mathbf{\mathcal{H}}}_{k},^{W}\mathcal{L}_{k},{{}^{W}\mathcal{M}_{k}}]\>k\in\mathcal{K}$ (30)
@@
regardless of the formulation. If the formulation used does not directly estimate a required output element, it must be computed directly from the available information. This behavior is enforced in DynoSAM’s back-end implementation via clear abstraction layers.
VI-D Implementation
@@
The camera pose and motion estimation algorithms described in the previous sections are implemented in C++. GTSAM 4.2 [54] is used for all non-linear optimization problems which are solved using Levenberg-Marquardt. For all experiments, unless otherwise specified (Sec. VII-E), full-batch optimization is used.
@@
At each time-step, the back-end component constructs a system by incorporating new factors and variables based on inputs from the front-end, which consists of initial estimates for the camera pose $\prescript{W}{}{\mathbf{X}}_{k}$, per-object motions $\prescript{W}{k-1}{\mathbf{\mathcal{H}}}_{k}$ and a set of visual measurements $\mathcal{Z}_{k}$. The system is then optimized, and the resulting estimates are processed to generate the output $\mathcal{O}_{k}$. Compared to a typical static SLAM system, the graph format of the input data is significantly more complex, and, depending on the formulation, may contain ternary Eq. (17) and quaternary edges Eq. (21).
@@
For both formulations discussed in Sec. V, the system defines new variables at each time-step for $\prescript{W}{k-1}{\mathbf{H}}_{k}$ or $\prescript{W}{}{\mathbf{L}}_{k}$ which represent the state of each object, and initializes them with values from the front-end where applicable. Each dynamic point measurement in $\mathcal{D}_{3D,k}$ contributes to constraining object states based on the observed camera pose(s) and its corresponding dynamic map point.
@@
This combination of factors and state variables results in a Dynamic SLAM problem that DynoSAM manages and constructs efficiently and flexibly, despite the significant increase in graph size and associated bookkeeping when compared to the classical static SLAM problem. DynoSAM achieves this by decoupling the measurement bookkeeping from the implementation of individual Dynamic SLAM formulations through abstraction layers, as mentioned in the previous section. This design choice dramatically simplifies the implementation of new formulations, allowing researchers to easily experiment with different approaches. Further technical details regarding graph construction and the abstractions used are available in our documented open-source software [^1] https://github.com/ACFR-RPG/DynOSAM.
@@
VII Experiments
Our framework has been rigorously tested on indoor and outdoor datasets featuring diverse dynamic objects. We present results of camera trajectories and object motions using full-batch optimization, as well as the performance using sliding window optimization (Sec. VII-E). In particular, we compare the accuracy of DynoSAM’s World-Centric Motion and Pose Estimator (henceforth referred to as DynoSAM Motion and DynoSAM Pose) against MVO [17], VDO-SLAM [15], ClusterSLAM [24] and DynaSLAM II [21]. Our evaluation is conducted using the metrics outlined in Sec. VII-A and provides a comprehensive assessment of these systems. All evaluation metrics are included in our open-source framework and are based on the implementation of evo [64]. Of the systems we compare against, only VDO-SLAM [15] provides an open-source implementation of their pipeline which we use to directly obtain the results reported. In the case of MVO, their pipeline’s raw output data was provided by the authors, allowing all metrics to be used during evaluation. ClusterSLAM [24] uniquely reports accumulated average camera and object pose errors over all sequences, instead of individual results per sequence. We therefore report an accumulated average error to facilitate comparison with their system and use the values reported in [24]. We use the camera pose errors of DynaSLAM II as reported in [21]. Due to limited reported results and the lack of object ID associations between their results and our dataset, we exclude DynaSLAM II object errors from our comparisons.
VII-A Metrics
TABLE I: Quantitative evaluations of camera trajectory for all experiments.

### v3

@@ V-B1 Feature Detection & Tracking
Static Feature Tracking. At each time-step, a sparse set of detected static keypoints $\mathcal{S}_{\text{2D},k}$ are tracked across consecutive frames using optical flow. Our framework offers options to select different feature extraction algorithms depending on the use case; for our experiments we used Shi-Tomasi corners [shi1994good] for indoor environments and ORB [rublee2011ORB] features for outdoor environments. By default, our front-end uses the Lucas-Kanade tracker [lucas1981iterative_LK] to generate feature correspondences. The input optical flow $I^{F}_{k}$ can also be used when it is reliable. Once tracked, our front-end performs geometric verification to discard poor correspondences.
@@
To improve robust tracking, we retain only relevant features by applying Adaptive Non-Maximal Suppression (ANMS) [bailo2018efficient_anms] to the detected keypoints. This procedure retains only the most informative and reliable features, while simultaneously promoting a spatially uniform distribution of features across the image. ANMS achieves this by culling uninformative features whilst ensuring that a minimum number of features remain ($800$ in our experiments). Our front-end will detect new keypoints on the static structure if the number of inlier tracks falls below this minimum threshold. This adaptive mechanism ensures that a consistently high number of well-distributed features are tracked across successive frames, thereby enhancing tracking robustness and accuracy.
@@
Dynamic Feature Tracking. We track a dense set of features $\mathcal{D}_{\text{2D},k}$ for each object. Dense tracking is important to achieve good coverage over the entire observable object for robust motion estimation [zhang20iros]. For each dynamic measurement we maintain an associated object-level label $j$, such that features can be associated to the same object between frames.
@@
The instance segmentation mask $I^{S}_{k}$ is used to retrieve the object label $j$ per pixel. While this mask can be generated from any standard instance segmentation network, many of these networks do not guarantee that the instance labels will be temporally consistent. Before tracking dynamic keypoints, object masks are tracked using [zhang2022bytetrack]. We then remap the original instance labels to a temporally consistent label $j$ per object. If the user has their own instance segmentation and tracking method, they can choose to bypass this algorithm. In that case, the user only needs to ensure that each pixel represents the unique tracking label $j$.
@@
Although the instance segmentation mask $I^{S}_{k}$ is able to separate objects from the background, it cannot distinguish between static and moving objects. To address this, we use the method of Zhang et al[zhang20iros] to identify moving objects based on scene flow, allowing us to focus on tracking features only on the dynamic objects. This module can be easily modified to integrate learning-based methods to identify moving objects.
@@
Once the moving objects in the image are identified and tracked, dense features are extracted by sampling keypoints uniformly within each tracked object mask. Optical flow $I^{F}_{k}$ is then used directly to find the correspondences between frames. Our tracking algorithm ensures that, where possible, a consistent and significant number of features ($800$ in our experiments) on each object are tracked. Similarly to static points, only inliers from the previous frame are used.
V-B2 Initial Camera Pose Estimation
@@
At each time-step, static measurements $\mathcal{S}_{k}$ are used to estimate the initial camera pose. As shown in the motion estimation module in Fig. 5, the first step is the PnP algorithm [ke17cvpr] which estimates $\prescript{W}{}{\mathbf{X}}_{k}$ by minimizing the re-projection error:
@@
between the tracked static keypoints in the current frame at time-step $k$ and the local map constructed from the previous frame. RANSAC verification and outlier rejection is performed using the implementation as provided by OpenGV [kneip2014opengv] to obtain robust initial estimation of the camera pose. This initial camera pose is further refined through joint optical flow and pose optimization method detailed in Sec. V-B4.
V-B3 Initial Object Motion Estimation
The initial object motion $\prescript{W}{k-1}{\mathbf{H}}_{k}\in\mathcal{H}_{k}$ is estimated following the method of Zhang et al[zhang2020vdoslam]:
@@ V-B3 Initial Object Motion Estimation
where $\prescript{W}{k-1}{\mathbf{G}}_{k}=\prescript{W}{}{\mathbf{X}}_{k}^{-1}\>\prescript{W}{k-1}{\mathbf{H}}_{k}$. Since (26) is in the same form as (24), we can solve for $\prescript{W}{k-1}{\mathbf{G}}_{k}$ in the same fashion, and the object motion:
@@
can be directly recovered. As before, RANSAC is used to detect outliers and update the inlier tracks in the current frame. Building on [zhang2020vdoslam], we enhance the initial object motion estimation and refine the inlier tracks through two additional steps, detailed in Sec. V-B4 and Sec. V-B5, respectively. The method proposed in  Sec. V-B4 was used in [zhang2020vdoslam] to improve the estimation and tracking. With the addition of  Sec. V-B5 our method augments this process to further improve the estimation and identify outliers. At each stage of the refinement process, the set of inlier tracks is updated, significantly improving the robustness of the estimation. The three-steps outlined assumes that the motion of each object is independent, allowing DynoSAM to parallelize the object motion estimation process by handling each object instance in a separate thread.
V-B4 Joint Optical-Flow Refinement
Following the work of [zhang2020vdoslam, zhang20iros], the optical-flow is jointly refined with the camera pose and object motions. This step ensures robust and accurate tracking of both static and dynamic features and helps to account for any errors in the initial optical-flow calculation. The measured flow and the associated $\mathrm{SE}(3)$ transformation are jointly refined by reformulating the re-projection errors from (26) and (24) in terms of the measured flow:
$\mathbf{r}_{f,H}$ $=\mathbf{z}^{i}_{\text{2D},k-1}+\mathbf{f}^{i}_{k-1,k}-\pi(\prescript{W}{k-1}{\mathbf{G}}_{k}\>\prescript{W}{}{\mathbf{m}}_{k-1}^{i})$ (27)
$\mathbf{r}_{f,X}$ $=\mathbf{z}^{i}_{\text{2D},k-1}+\mathbf{f}^{i}_{k-1,k}-\pi(\prescript{W}{}{\mathbf{X}}_{k}^{-1}\>\prescript{W}{}{\mathbf{m}}^{i})$
@@ V-B4 Joint Optical-Flow Refinement
where $\theta$ represents either $\prescript{W}{}{\mathbf{X}}_{k}$ camera pose or $\prescript{W}{k-1}{\mathbf{G}}_{k}$ in the case of object motion estimation. Depending on $\theta$, $\mathbf{r}_{f,\theta}$ is either $\mathbf{r}_{f,X}$ or $\mathbf{r}_{f,H}$, as in (27). The covariance matrix $\Sigma_{0}\in\mathbb{R}^{{2\times 2}}$ associated with the flow prior $\mathbf{r}_{0}(f)$ is diagonal and associated with the measured optical-flow. As each object moves independently, we construct and solve this optimization problem in parallel per object. During optimization, we enhance robustness by eliminating additional outliers, specifically those exhibiting large re-projection errors.
V-B5 Object Motion Refinement
The object motion estimates can be further refined by employing the rigid-body motion model defined in (16) to formulate an additional nonlinear least-square optimization problem that directly estimates the object motion $\prescript{W}{k-1}{\mathbf{H}}_{k}$. In contrast to the preceding step, which uses a 2D pixel error, this residual function is based on 3D point errors. Empirically, we have observed that this 3D-based approach significantly enhances estimation accuracy and removes additional outliers not previously identified. Using the motion estimate from the previous step as an initial estimate, the motion only refinement graph shown in Fig. 6 is build per object and defines the nonlinear least-squares problem:
@@ V-B5 Object Motion Refinement
The Levenberg–Marquardt solver is used to obtain the final object motion estimate $\prescript{W}{k-1}{\mathbf{H}}_{k}$. Based on (24), (29b) forms projection factors with an associated covariance matrix $\Sigma_{2D,R}\in\mathbb{R}^{2\times 2}$, connecting dynamic points and the observing camera poses at time-step k1k-1 and $k$. (29c) describes the 3D motion residual of (16) and connects the common motion with each tracked landmark. $\Sigma_{3D,H}\in\mathbb{R}^{3\times 3}$ is the associated covariance matrix. Equation (29a) represents prior factors on the camera poses. Since this refinement step focuses exclusively on improving object motion estimation, and not camera pose, a strong prior on the camera poses are used during optimization. This is achieved by assigning the covariance matrix $\Sigma_{0}\in\mathbb{R}^{6\times 6}$ with very small values ($\sigma=0.0001$ in our experiments).
@@
Figure 6: Factor graph representing the motion only refinement graph. For each object $j\in\mathcal{J}_{k}$, the motion is refined by minimizing the 3D motion residual using measurements of points tracked between k1k-1 and $k$. In our example we show the dynamic points in orange, the motion in green, the 3D motion residual in purple and the re-projection factors in white. We only show two sets of points $i$ and $i+n$, while in reality there would be $n$ motion factors.
Figure 7: Example scenarios from datasets used: (a): KITTI tracking, (b): OMD, (c): Outdoor Cluster, (d): TartanAir (Shibuya), (d): VIODE.
V-C Back-end
@@
The back-end fuses static and dynamic measurements to jointly estimate the trajectory of the camera and each object as well as a global map of static and dynamic points. Our back-end allows different Dynamic SLAM formulations to be selected and used for estimation. Full-batch and sliding window methods are provided for optimisation.
@@
As previously discussed, formulations can represent objects using different state variables (poses or motions as shown in this work) and in different frames of reference (locally as in DynaSLAM II [bescos2021ral] or in the camera frame as in MVO [judd2024ijrr_mvo]). Therefore, to support modularity, evaluation and use in downstream applications DynoSAM defines a common interface between the formulation the system output. A pose-estimation processing step is implemented that converts all estimated states into a common output:
$\mathcal{O}_{k}=[\prescript{W}{}{\mathbf{X}}_{k},\prescript{W}{k-1}{\mathbf{\mathcal{H}}}_{k},{{}^{W}\mathcal{L}_{k}},{{}^{W}\mathcal{M}_{k}}]\>k\in\mathcal{K}$ (30)
@@
If the formulation used does not directly estimate a required output element, it must be computed directly from the available information. This behavior is enforced in DynoSAM’s back-end implementation via clear layers of abstraction.
V-D Implementation
@@
The camera pose and motion estimation algorithms described in the previous sections are implemented in C++. GTSAM 4.2 [gtsam] is used for all non-linear optimization problems which are solved using Levenberg-Marquardt. For all experiments, unless otherwise specified (Sec. VI-E), full-batch optimization is used.
@@
At each time-step, the back-end component constructs a system by incorporating new factors and variables based on inputs from the front-end, which consists of initial estimates for the camera pose $\prescript{W}{}{\mathbf{X}}_{k}$, per-object motions $\prescript{W}{k-1}{\mathbf{\mathcal{H}}}_{k}$ and a set of visual measurements $\mathcal{Z}_{k}$. After optimisation, the estimates are processed to generate the output $\mathcal{O}_{k}$.
@@
For both formulations discussed in Sec. IV, the system defines new variables at each time-step for $\prescript{W}{k-1}{\mathbf{H}}_{k}$ or $\prescript{W}{}{\mathbf{L}}_{k}$ which represent the state of each object, and initializes from the front-end where applicable.
@@
The combination of factors and state variables results in a Dynamic SLAM problem that, compared to a typical static SLAM system, is significantly more complex, and may contain ternary (16) and quaternary edges (21). Despite the significant increase in graph size and associated bookkeeping DynoSAM efficiently manages the construction of each graph. DynoSAM achieves this by decoupling the measurement bookkeeping from the implementation of individual Dynamic SLAM formulations through abstraction layers, as mentioned in the previous section. This design choice dramatically simplifies the implementation of new formulations, allowing researchers to easily experiment with different approaches. Further technical details regarding graph construction and the abstractions used are available in our documented open-source software.
@@
VI Experiments
The primary focus of this work is the estimation object motion and pose. Hence, we evaluate both DynoSAM’s WCME (Sec. IV-B) and WCPE (Sec. IV-C) formulations on a range of indoor and outdoor datasets featuring diverse dynamic objects and contain ground truth for both camera pose and object motion/pose. To facilitate a wide range of comparison we evaluate DynoSAM on the KITTI tracking [Geiger13ijrr], OMD [Judd19ral] and Outdoor Cluster [Huang2019iccv] [^2] The datasets are provided in addition to the paper and can be found at https://huangjh-pub.github.io/page/clusterslam-dataset/. Note that the ground truth rotation is actually in $[qw,qx,qy,qz]$ form instead of the $[qx,qy,qz,qw]$ format stated on the website. We use VDO-SLAM [zhang2020vdoslam], MVO [judd2024ijrr_mvo] and ClusterSLAM [Huang2019iccv] as state-of-the-art Dynamic SLAM systems to assess our object motion/pose estimation. Furthermore, we propose a new evaluation metric for object motion and discuss how it better facilitates comparison across systems.
In addition to object motion/pose, we evaluate our camera pose estimation by comparing against DynaSLAM II [bescos2021ral], VDO-SLAM [zhang2020vdoslam], MVO [judd2024ijrr_mvo], ClusterSLAM [Huang2019iccv], AirDOS [Qiu2022icra_airdos], ORB-SLAM3 [Campos2021tro] and DynaVINS [Song2022ral_dynavins]. Datasets that do not contain ground truth object information, such as TartanAir Shibuya [Qiu2022icra_airdos] and VIODE [minoda2021viode], are also included to validate the accuracy and robustness of our camera pose estimation in highly dynamic and challenging environments. Our evaluation is conducted using the metrics outlined in Sec. VI-B and provides a comprehensive assessment of these systems. All evaluation metrics are included in our open-source framework and are based on the implementation of evo [grupp2017evo].
VI-A Datasets and Third Party Results
VI-A1 Sequences with Camera & Object Ground Truth
We use the KITTI tracking dataset [Geiger13ijrr], Outdoor Cluster [Huang2019iccv] and OMD [Judd19ral] (Fig. 7 a-c) to evaluate object motion and pose estimation in scenes with multiple objects, complex motion, occlusion, and semi-rigid entities such as cyclists. We use a processed version of the KITTI tracking dataset containing ground truth information per object per frame. Outdoor Cluster is simulated and provides ground truth. OMD uses a Vicon system to capture all ground truth trajectories [Judd19ral].
KITTI and Outdoor Cluster represent large-scale outdoor driving scenarios with numerous moving objects. However, most of these objects experience relatively constant motion. We therefore additionally include swinging_4_unconstrained (S4U), the OMD sequence with the most challenging dynamics, which features $4$ swinging cubes with unconstrained and unpredictable motions. The hand-held camera motion further increases the difficulty. While MVO uses another OMD sequence to test their object re-association ability after occlusion, this functionality is not relevant for motion evaluation and is therefore not assessed in this work.
VI-A2 Challenging Sequences with Only Camera Ground Truth
While not our primary focus, evaluating camera pose estimation is crucial to demonstrate DynoSAM’s accuracy and robustness in challenging conditions, including dynamic objects, occlusions, and low light. We test on the simulated TartanAir Shibuya [Qiu2022icra_airdos] and VIODE [minoda2021viode] datasets. TartanAir features over $30$ moving humans with complex motions that frequently occlude the view (Fig. 7 (d)). VIODE poses further challenges with full occlusion scenarios as shown in Fig. 7 (e), low-light, and includes UAV-like motion with extreme translation and rotation. This dataset additionally contains IMU.
VI-A3 Obtaining Third Party Results For Object Motions
Obtaining consistent object motion/pose evaluation from third-party systems is challenging due to limited open-source options and different metrics and formats. Of the Dynamic SLAM systems we compare against, only VDO-SLAM [zhang2020vdoslam] is open-source, and we used it to generate results directly. To ensure fairness we run VDO-SLAM on the same processed data as DynoSAM. VDO-SLAM shares the underlying motion formulation with WCME but differs in its use of g2o [Kummerle11icra] for optimizations, as well as in its front-end design and tracking strategy. MVO [judd2024ijrr_mvo] reports object motions for a few sequences, and the authors kindly provided results on S4U and KITTI $00$ for our evaluation. ClusterSLAM [Huang2019iccv] uniquely reports average camera and object pose errors accumulated over all Outdoor Cluster sequences. We therefore report an accumulated average error on this dataset and compare against the values reported by Huang et al[Huang2019iccv]. DynaSLAM II [bescos2021ral] provides per-object pose errors for a few selected objects per sequence. On the KITTI dataset, without source code, we cannot align object IDs and thus unfortunately compare only camera pose for these sequences reported in their paper. We however note that DynoSAM improves upon DynaSLAM II’s reported object pose estimation by $91\%$ in rotation and $36\%$ in translation.
VI-B Metrics
This section details the metrics used in our DynoSAM framework to evaluate the accuracy of pose and motion estimations for camera and objects. For any error $\mathbf{E}_{k}\in\mathrm{SE}(3)$ at time-step $k$, we compute and report the root-mean-squared error (RMSE) of the translation and rotation component separately throughout a sequence:
$\mathbf{E}_{t}$ $=\text{RMSE}(trans\left(\mathbf{E_{k}}\right))$ (31)
$\mathbf{E}_{r}$ $=\text{RMSE}(rot\left(\mathbf{E_{k}}\right))$
$\text{RMSE}(e)$ $=\sqrt{\frac{1}{n(\mathcal{K})}\sum_{k\in\mathcal{K}}\lVert e_{k}\rVert^{2}}\text{,}$
where $e_{k}$ is the scalar error of either the $L_{2}$ norm of the translational component, denoted as $trans\left(\mathbf{E_{k}}\right)$, or the angle of the rotational component, denoted as $rot\left(\mathbf{E_{k}}\right)$, of $\mathbf{E}_{k}\in\mathrm{SE}(3)$.
VI-B1 Pose Evaluation
Camera and object pose estimation accuracy is evaluated using Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) as defined by Sturm et al[sturm2012iros]. Given a ground truth transformation $\mathbf{M}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}, k}}\in\mathrm{SE}(3)$} at $k$ and a corresponding reference estimate $\mathbf{M}_{\text{k}}\in\mathrm{SE}(3)$, the ATE is defined as:
$\mathbf{ATE}=\text{RMSE}\big(trans\left(\mathbf{M}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}^{-1}\>\mathbf{M}_{k}\right)\big)\text{.}$ (32)
The ground truth reference frame is emphasized in red to clearly distinguish it from the estimated reference frame.
ATE is commonly used to assess the global consistency in trajectory estimation [mur2017orb, bescos2021ral, huang2020cvpr] and we use ATE to evaluate the accuracy in camera pose. We do not evaluate ATE for objects since it greatly depends on the ATE of the camera and how the object frame is defined for each system.
TABLE I: Quantitative evaluations of camera trajectory (ATE and RPE) against other Dynamic SLAM systems. Entries marked with ‘-’ indicate that results are not available for that sequence. Best results are marked in bold and second best with underscore.